OpenStudy (anonymous):

Proving the following trig identities: sin^2 (x/2)= (csc x - cot x)/ (2 csc x) help please

OpenStudy (anonymous):

i get that (1-cosx)/2 = (csc x - cot x)/ (2 csc x) does it help ?

OpenStudy (anonymous):

you need to find the relation between (1-cosx)/2 and sin^2 (x/2)=

OpenStudy (anonymous):

thank you for explaining, this helps a lot :)

OpenStudy (anonymous):

you didn't forgot any square in sin^2 (x/2)= (csc x - cot x)/ (2 csc x) ?

OpenStudy (anonymous):

thank you i proved it with your help :)

OpenStudy (anonymous):

how did you do it :o ?

OpenStudy (anonymous):

The LHS is just (1/2) (1+ cos(x)), so it becomes cos(2a) = 1 - sin^2(a) (standard) with a = x/2

OpenStudy (anonymous):

1-cos(x) *** My bad

OpenStudy (anonymous):

I still don't get the last part :(